OpenStudy (anonymous):

what is the derivative of x^(1/x)?? @MIT 18.02 Multiva…

OpenStudy (jamesj):

Let y = x^(1/x). Then ln y = 1/x . ln x hence differentiating wrt x 1/y . y' = 1/x . 1/x + -1/x^2 . ln x Hence y' = 1/x^2 (1 - ln x) y = 1/x^2 (1 - ln x) x^(1/x)

OpenStudy (jamesj):

I.e., \[ \frac{d\ }{dx} x^{1/x} = \frac{1 - \ln x}{x^2} x^{1/x} \]

OpenStudy (anonymous):

\[y=x^{\frac{1}{x}}\Longrightarrow\ln y = \frac{1}{x}\ln x\Longrightarrow \frac{1}{y}\cdot\frac{dy}{dx}=\left(-\frac{1}{x^2}\right)\ln x+\frac{1}{x}\cdot\frac{1}{x}\]\[\Longrightarrow\frac{dy}{dx}=y\left(\frac{1-\ln x}{x^2}\right)=x^{\frac{1}{x}}\left(\frac{1-\ln x}{x^2}\right)\]

OpenStudy (anonymous):

ooo ok thx